# Nesterov acceleration in benignly nonconvex landscapes

### v1

@@ B.2. Proof of Theorem 11: Acceleration in Discrete Time (Deterministic setting)
(Gupta et al., 2023, Lemma 13) Assume that $\nabla f$ is Lipschitz-continuous with Lipschitz-constant $L$. Then
@@
Setup. Denote by $P^{\bot}=I-P$ the orthogonal projection onto the orthogonal complement of the space $V$ which $P$ projects onto. Consider the Lyapunov sequence defined by
${\mathcal{L}}_{n}=f(x_{n})-\inf f+\frac{1}{2}\|\ P^{\bot}v_{n}+x_{n}^{\prime}-\pi(x_{n}^{\prime})\|^{2}+\frac{(1+\sqrt{\mu\eta})^{2}}{2(1-\sqrt{\mu\eta})}\|Pv_{n}\|^{2}.$
@@
$\|\ P^{\bot}v+x-\pi(x)\|^{2}+\frac{(1+\sqrt{\mu\eta})^{2}}{2(1-\sqrt{\mu\eta})}\|Pv\|^{2}\to\|\ P^{\bot}v+x-\pi(x)\|^{2}+\|Pv\|^{2}=\|v+x-\pi(x)\|^{2}$
@@
Step 2. Denote $P^{\bot}=I-P$ the orthogonal projection onto the orthogonal complement of the space that $P$ projects onto. We compute
@@
$=\frac{1-\sqrt{\mu\eta}}{2}\left(\|Pv_{n}\|^{2}+\eta\|P\nabla f(x^{\prime}_{n})\|^{2}-2\sqrt{\eta}\langle\nabla f(x^{\prime}_{n}),Pv_{n}\rangle\right)$
@@
$L_{n+1}$ $=\frac{(1-\sqrt{\mu\eta})^{2}}{2}\|P^{\bot}v_{n}\|^{2}+\frac{\mu}{2}\|P^{\bot}x^{\prime}_{n}\|^{2}+\frac{\eta}{2}\|P^{\bot}\nabla f(x^{\prime}_{n})\|^{2}$
@@
$L_{n+1}$ $\leq\frac{(1-\sqrt{\mu\eta})^{2}}{2}\|P^{\bot}v_{n}\|^{2}+\frac{\mu}{2}\|P^{\bot}x^{\prime}_{n}\|^{2}+\frac{\eta}{2}\|P^{\bot}\nabla f(x^{\prime}_{n})\|^{2}$
@@
$L_{n+1}$ $\leq\frac{(1-\sqrt{\mu\eta})^{2}+(1-\sqrt{\mu\eta})\eta\varepsilon}{2}\|P^{\bot}v_{n}\|^{2}+\frac{(1-\sqrt{\mu\eta})(1+\eta\varepsilon)}{2}\|Pv_{n}\|^{2}$
@@
$L_{n+1}$ $\leq(1-\sqrt{\mu\eta})(f(x_{n})-\inf f)+\frac{1-\sqrt{\mu\eta}}{2}\|P^{\bot}v_{n}\|^{2}$
@@
B.3. Proof of Theorems 13 and 14: Acceleration with additive noise
@@
A proof can be found in (Gupta et al., 2023, Lemma 15). The second identity, which is not included, can be analogously. As an application, we prove a stochastic analogue of Lemma 19.
@@
Finally, we provide an auxiliary result to resolve a slightly more complicated recursion.
@@
${\mathbb{E}}[f(x_{n+1}]\leq{\mathbb{E}}[f(x_{n}^{\prime})]-\frac{\eta}{2}\,\|\nabla f(x_{n}^{\prime})\|^{2}+\frac{L\sigma_{a}^{2}\eta^{2}}{2}$
by Lemma 21. In the quadratic term, we need to take the expectation of $g_{n}$ rather than $f(x_{n}^{\prime})$. By construction, the terms involving ${\mathbb{E}}[\|\nabla f(x_{n}^{\prime})\|^{2}]$ still balance with the same parameters, leading to an additional contribution of $\eta\,\sigma_{a}^{2}$ due to the stochastic gradient estimates. Since $\eta\leq 1/L$, we can bound $L\eta^{2}\leq\eta$ so overall we obtain the estimate
@@
B.4. Proof of Theorem 15: Acceleration with both additive and multiplicative noise
In the convex setting, Gupta et al. (2023) state a version of Theorem 15 in slightly greater generality in terms of choosing variables. The same more general proof goes through also here.
@@ B.4. Proof of Theorem 15: Acceleration with both additive and multiplicative noise
We note that the coefficient of the norm of the tangential gradient is $(b-\sqrt{\mu\alpha})\sqrt{\alpha}\,\gamma\leq\gamma^{2}$ by the definition of $\gamma=(b-\sqrt{\mu\alpha})\sqrt{\alpha}+\sqrt{\mu}\eta$.
Next, we combine this estimate with the bound on $f(x_{n+1})$ from Step 1 and use the geometric conditions (6) and (7) on $f$ to control the inner products of $\nabla f(x_{n}^{\prime})$ with $v_{n}$ and $P^{\bot}x_{n}^{\prime}$ in the previous expression as well as the variance bound (9) for the gradient estimates:
@@
$\quad+\frac{\mu}{2}\,{\mathbb{E}}\big{[}\|P^{\bot}x_{n}^{\prime}\|^{2}\big{]}+\frac{\lambda}{\,}{\mathbb{E}}\big{[}\|Pv_{n}\|^{2}\big{]}\bigg{\}}+\frac{L\eta^{2}+\frac{\eta}{1+\sigma_{m}^{2}}}{2}\,\sigma_{a}^{2}$
@@ Appendix C A brief comparison of geometric conditions for optimization
C.1. Definitions and elementary properties
@@
  5. (5)
    satisfies the first order $\mu$-strong convexity condition with respect to the closest minimizer if for all $x$ we have
@@
  1. (1)
    If $f$ is first order $\mu$-strongly convex and has a minimizer in $U$, then $f$ satisfies the first order $\mu$-strong convexity condition with respect to the closest minimizer.
  2. (2)
    If $f$ satisfies the first order $\mu$-strong convexity condition with respect to the closest minimizer, then $f$ satisfies the PL condition with the same constant $\mu$.
  3. (3)
    If $f$ is $\mu$-strongly convex with respect to the closest minimizer on $\mathcal{U}_{\alpha}=\{x\in\mathbb{R}^{d}:f(x)<\alpha\}$, then the line segment connecting $x$ and $\pi(x)$ is contained in $\mathcal{U}_{\alpha}$ for all $x\in\mathcal{U}_{\alpha}$.
  4. (4)
@@
In particular, for every $x$ there exists a unique closest minimizer $\pi(x)=x^{*}$. The first order convexity condition with respect to the closest minimizer therefore requires the first order convexity condition only for pairs of points $x,x^{*}$ rather than all points $x,z$.
Second claim. This result follows by the same proof that implies the PL condition for strongly convex functions: If $f$ satisfies a first order $\mu$-strong convexity condition with respect to the closest minimizer, then
@@ C.2. A one-dimensional example
First order strong convexity with respect to the closest minimizer. Note that
@@
In particular, $f$ is $\mu$-strongly convex with respect to the global minimizer with
@@
if $\varepsilon\sqrt{1+4R^{2}}<1$ and it fails to be $\mu$-strongly convex with respect to the global minimizer for any $\mu>0$ otherwise.
Quasar-convexity. Since the minimizer is unique, $(\gamma,\mu)$-strong convexity is a strictly more general notion than first order strong convexity with respect to the closest minimizer as we have an additional parameter $\gamma$ to relax the requirements. Essentially the same calculation reads
@@
PL condition $\mu$-sc wrt minimizer $\mu$-strongly convex
@@
Evidently, classical strong convexity implies strong convexity with respect to the global minimizer which in turn implies the PL condition. The parameter ranges and constants are generally vastly different. All estimates, except for the PL constant, are sharp.
In particular, the one-dimensional examples demonstrate that first order convexity with respect to the closest minimizer is strictly weaker than convexity along the line segment connecting $x$ to $\pi(x)$.
C.3. Deep learning
@@
While there are guarantees that individual algorithms escape certain types of critical points almost surely (e.g. strict saddles, (Lee et al., 2019; O’Neill & Wright, 2019)), they may take very long to do so (Du et al., 2017). The analysis of accelerated rates becomes asymptotic at best. We claim that our notion of strong convexity with respect to the closest minimizer suffers from the same ‘optimism’ globally, but locally captures two important features of deep learning landscapes close to the set of global minimizers which are not captured by concepts which require a geometric condition with respect to all minimizers: A manifold along which we can move tangentially, and convexity in directions which are perpendicular to the manifold.

### v2

@@ B.2 Proof of Theorem 11: Acceleration in Discrete Time (Deterministic setting)
(Gupta et al., 2024, Lemma 13) Assume that $\nabla f$ is Lipschitz-continuous with Lipschitz-constant $L$. Then
@@
Setup. Denote by $P^{\bot}=I-P$ the orthogonal projection onto the orthogonal complement of the space which $P$ projects onto. Consider the Lyapunov sequence defined by
${\mathcal{L}}_{n}=f(x_{n})-\inf f+\frac{1}{2}\|\ P^{\bot}v_{n}+\sqrt{\mu}(x_{n}^{\prime}-\pi(x_{n}^{\prime}))\|^{2}+\frac{(1+\sqrt{\mu\eta})^{2}}{2(1-\sqrt{\mu\eta})}\|Pv_{n}\|^{2}.$
@@
$\|\ P^{\bot}v+\sqrt{\mu}(x-\pi(x))\|^{2}+\frac{(1+\sqrt{\mu\eta})^{2}}{2(1-\sqrt{\mu\eta})}\|Pv\|^{2}\to\|\ P^{\bot}v+\sqrt{\mu}(x-\pi(x))\|^{2}+\|Pv\|^{2}=\|v+\sqrt{\mu}(x-\pi(x))\|^{2}$
@@
Step 2. We compute
@@
$=\frac{1-\sqrt{\mu\eta}}{2}\left(\|Pv_{n}\|^{2}+\eta\|P\nabla f(x^{\prime}_{n})\|^{2}-2\sqrt{\eta}\langle\nabla f(x^{\prime}_{n}),Pv_{n}\rangle\right).$
@@
${\mathcal{L}}_{n+1}$ $=\frac{(1-\sqrt{\mu\eta})^{2}}{2}\|P^{\bot}v_{n}\|^{2}+\frac{\mu}{2}\|P^{\bot}x^{\prime}_{n}\|^{2}+\frac{\eta}{2}\|P^{\bot}\nabla f(x^{\prime}_{n})\|^{2}$
@@
${\mathcal{L}}_{n+1}$ $\leq\frac{(1-\sqrt{\mu\eta})^{2}}{2}\|P^{\bot}v_{n}\|^{2}+\frac{\mu}{2}\|P^{\bot}x^{\prime}_{n}\|^{2}+\frac{\eta}{2}\|P^{\bot}\nabla f(x^{\prime}_{n})\|^{2}$
@@
${\mathcal{L}}_{n+1}$ $\leq\frac{(1-\sqrt{\mu\eta})^{2}+(1-\sqrt{\mu\eta})\eta\varepsilon}{2}\|P^{\bot}v_{n}\|^{2}+\frac{(1-\sqrt{\mu\eta})(1+\eta\varepsilon)}{2}\|Pv_{n}\|^{2}$
@@
${\mathcal{L}}_{n+1}$ $\leq(1-\sqrt{\mu\eta})(f(x_{n})-\inf f)+\frac{1-\sqrt{\mu\eta}}{2}\|P^{\bot}v_{n}\|^{2}$
@@
B.3 Proof of Theorems 13 and 14: Acceleration with additive noise
@@
A proof can be found in (Gupta et al., 2024, Lemma 15). The second identity, which is not included therein, can be proved analogously. As an application, we prove a stochastic analogue of Lemma 19.
@@
Finally, we provide an auxiliary result to resolve a recursion.
@@
${\mathbb{E}}[f(x_{n+1}]\leq{\mathbb{E}}[f(x_{n}^{\prime})]-\frac{\eta}{2}\,{\mathbb{E}}\big{[}\|\nabla f(x_{n}^{\prime})\|^{2}\big{]}+\frac{L\sigma_{a}^{2}\eta^{2}}{2}$
by Lemma 21. In the quadratic terms in steps 2 and 3, we need to take the expectation of $g_{n}$ rather than $\nabla f(x_{n}^{\prime})$. By construction, the terms involving ${\mathbb{E}}[\|\nabla f(x_{n}^{\prime})\|^{2}]$ still balance with the same parameters, leading to an additional contribution of $\eta\,\sigma_{a}^{2}$ due to the stochastic gradient estimates. Since $\eta\leq 1/L$, we can bound $L\eta^{2}\leq\eta$ so overall we obtain the estimate
@@
B.4 Proof of Theorem 15: Acceleration with both additive and multiplicative noise
In the strongly convex setting, Gupta et al. (2024) state a version of Theorem 15 in slightly greater generality in terms of choosing variables. The same more general proof goes through also here after we account for the tangential and normal components of the velocity as in the proof of Theorem 11.
@@ B.4 Proof of Theorem 15: Acceleration with both additive and multiplicative noise
We note that the coefficient of the norm of the tangential gradient is $(b-\sqrt{\mu\alpha})\sqrt{\alpha}\,\gamma\leq\gamma^{2}$ by the definition of $\gamma=(b-\sqrt{\mu\alpha})\sqrt{\alpha}+\sqrt{\mu}\eta$. Next, we combine this estimate with the bound on $f(x_{n+1})$ from Step 1 and use the geometric conditions (6) and (7) on $f$ to control the inner products of $\nabla f(x_{n}^{\prime})$ with $v_{n}$ and $P^{\bot}x_{n}^{\prime}$ in the previous expression as well as the variance bound (9) for the gradient estimates:
@@
$\quad+\frac{\mu}{2}\,{\mathbb{E}}\big{[}\|P^{\bot}x_{n}^{\prime}\|^{2}\big{]}+\frac{\lambda}{2}\,{\mathbb{E}}\big{[}\|Pv_{n}\|^{2}\big{]}\bigg{\}}+\frac{L\eta^{2}+\frac{\eta}{1+\sigma_{m}^{2}}}{2}\,\sigma_{a}^{2}$
@@ Appendix C A brief comparison of geometric conditions for optimization
C.1 Definitions and elementary properties
@@
  5. 5.
    satisfies the first order $\mu$-strong aiming condition if for all $x$ we have
@@
  1. 1.
    If $f$ is first order $\mu$-strongly convex and has a minimizer in $U$, then $f$ satisfies the $\mu$-strong aiming condition.
  2. 2.
    If $f$ satisfies the $\mu$-strong aiming condition, then $f$ satisfies the PL condition with the same constant $\mu$.
  3. 3.
    If $f$ is $\mu$-strongly aiming on $\mathcal{U}_{\alpha}=\{x\in\mathbb{R}^{d}:f(x)<\alpha\}$, then the line segment connecting $x$ and $\pi(x)$ is contained in $\mathcal{U}_{\alpha}$ for all $x\in\mathcal{U}_{\alpha}$.
  4. 4.
@@
In particular, for every $x$ there exists a unique closest minimizer $\pi(x)=x^{*}$. The strong aiming condition therefore requires the first order convexity condition only for pairs of points $x,x^{*}$ rather than all points $x,z$.
Second claim. This result follows by the same proof that implies the PL condition for strongly convex functions: If $f$ satisfies the $\mu$-strong aiming condition, then
@@ C.2 A one-dimensional example
$\mu$-strong aiming condition. Note that
@@
In particular, $f$ is $\mu$-strongly aiming (with respect to the unique global minimizer) with
@@
if $\varepsilon\sqrt{1+4R^{2}}<1$ and it fails to be $\mu$-strongly aiming for any $\mu>0$ otherwise.
Quasar-convexity. Since the minimizer is unique, $(\gamma,\mu)$-strong convexity is a strictly more general notion than strong aiming condition as we have an additional parameter $\gamma$ to relax the requirements. Essentially the same calculation reads
@@
PL condition $\mu$-strongly aiming $\mu$-strongly convex
@@
Evidently, classical strong convexity implies strong aiming condition with respect to the global minimizer which in turn implies the PL condition. The parameter ranges and constants are generally vastly different. All estimates, except for the PL constant, are sharp. In particular, the one-dimensional examples demonstrate that strong aiming condition is strictly weaker than convexity along the line segment connecting $x$ to $\pi(x)$.
Figure 4: We compare the trajectories of gradient descent and Nesterov’s algorithm for the objective function $f$ in Example 25 with $R=6$ and $\varepsilon=0.075$ (left), $\varepsilon=0.08$ (middle) and $\varepsilon=0.085$ (right). Evidently, if $\varepsilon\sqrt{1+4R^{2}}$ is very close to the threshold value 1, gradient descent outperforms Nesterov’s algorithm with the theoretically guaranteed parameters.
In the case of quadratic objective functions or functions modelled on them (such as the distance function from a manifold), the PL constant is essentially the same as the parameter of strong convexity. For the function $f$ in Example 25 on the other hand, the PL constant is noticeably larger than the parameter of strong convexity with respect to the unique minimizer. In Figure 4, we observe that with the parameters $\eta=1/L$ and $\rho=(1-\sqrt{\mu\eta})/(1+\sqrt{\mu\eta})$, gradient descent may at times converge faster, at least with the parameter choice that is derived over a large function class rather than for an individual objective function.
C.3 Deep learning
@@
While there are guarantees that individual algorithms escape certain types of critical points almost surely (e.g. strict saddles, (Lee et al., 2019; O’Neill & Wright, 2019)), they may take very long to do so (Du et al., 2017). The analysis of accelerated rates becomes asymptotic at best. We claim that our notion of strong aiming condition suffers from the same ‘optimism’ globally, but locally captures two important features of deep learning landscapes close to the set of global minimizers which are not captured by concepts which require a geometric condition with respect to all minimizers: A manifold along which we can move tangentially, and convexity in directions which are perpendicular to the manifold.